4.0 mm 0 10 nC 3.0 mm

What is the proton's speed as it passes through point P?
Answer #1

Solution)

Let its initial speed be vi = 1.8 *10^6 m/s

Given,

d1 = 3 mm = 0.003 m

d2 = 4 mm = 0.004 m

let final speed be vf

Now,

We know, decrease in kinetic energy = increase in potential energy

So,

0.5 * 1.67 * 10^-27 * ((1.8 *10^6)^2 - vf^2) = 9*10^9 * 10 *10^-9 * 1.602 *10^-19 * (1/.003 - 1/.004)

We get,

vf = 1.342 *10^6 m/s (Ans)

=======

Good luck!:)